h
7. Let A = 134 -1 2 6 6 0 6 3 9 0 9 2 -3 -3 0 Find basis for Nul(A) and Col(A). 3 و 3
Answer #1

7) Given 2 3 4 0 -3 Az 6 3 9 3 - 6 3 9 0...9 basis of Null (A). We need to solve w 4 21 -3 0 2 6 6 O 39 3 6 Ry → Ry-3R, [3 0if we take the variables as b d 1 (ol 0 0 0 2 ک 10 U OO 9 o d. 0 0 0 12 6 at3b + 4c-dte =0 9+3 6+4c=0 a = -3b-ue 2d se = 0Basis for collari ف) 4 -1 ! O Ru Ru-682 ک۔ 2 6 o 0 0 0 0 9 lo 10 6 D12 o O 3 ilo 0 0 2 و اک۔ R₂ R3, O q ol ។ O 0 0 0 361 D Ru